# Microscopic Mechanism of Anyon Superconductivity Emerging from Fractional Chern Insulators

### v1

@@ Appendix E Counting for edge spectrum
We outline how the edge spectrum and its degeneracies can be obtained for a fixed charge sector $Q$, focusing on $m=n-1$ relevant for the $(\bar{1}\bar{1}2)$ state. We first discuss $Q=0$, where there is a unique ground state with $L_{c}=L_{s}=0$ and momentum $\mathcal{P}=0$. There are three dispersing modes: A pure charge mode with $L_{c}\neq 0$, $L_{s}=0$, a pure spin mode with $L_{s}\neq 0$, $L_{c}=0$, and a mixed mode with $L_{c}\neq 0\neq L_{s}$. For the charge mode, $L_{s}=0$ fixes $S_{z}=0$. The spectrum is given by
@@
resulting in the (1,1,2,3,..)(1,1,2,3,..)( 1 , 1 , 2 , 3 , . . ) counting of a single bosonic chiral mode. The counting for the spin mode is more involved. We have
@@ Appendix F Different cylinder geometries
[…]
Appendix G Quantum critical point in terms of Cooper pairs
@@
where $\alpha$ and $\tilde{\alpha}$ are emergent gauge fields describing the bosonic hierarchy state, and $\varphi$ is a scalar that drives the crystallization transition. The plateau transition between the $2/3$ hierarchy state of Cooper pairs (corresponding to the fermionic FCI) and the $1/2$ parent state (corresponding to the semion crystal) is now described by a condensation transition of the $\varphi$ boson. If $\varphi$ is condensed (Mott) we obtain the $\nu^{\text{Cooper}}=2/3$ ($\nu^{\text{Cooper}}=1/2$) state of Cooper pairs. From the equations of motion we notice that the total electron density is given by $2\mathrm{d}\alpha/2\pi$, the density of $\varphi$ bosons is $\mathrm{d}\tilde{\alpha}/2\pi$ and the flux seen by the bosons $\mathrm{d}\beta=-3\mathrm{d}\alpha=-3\mathrm{d}\tilde{\alpha}/2$. The superconductor is then obtained when the bosons form a FQH state corresponding to their filling of $-2/3$.

### v2

@@ Appendix E Counting for edge spectrum
We outline how the edge spectrum and its degeneracies can be obtained for a fixed charge sector $Q$, focusing on m=n1m=n-1 relevant for the $(\bar{1}\bar{1}2)$ state. We first discuss $Q=0$, where there is a unique ground state with $L_{c}=L_{s}=0$ and momentum $\mathcal{P}=0$. There are three dispersing modes: A pure charge mode with $L_{c}\neq 0$, $L_{s}=0$, a pure spin mode with $L_{s}\neq 0$, $L_{c}=0$, and a mixed mode with $L_{c}\neq 0\neq L_{s}$. For the charge mode, $L_{s}=0$ fixes $S_{z}=0$. The spectrum is given by
@@
resulting in the $(1,1,2,3,..)$ counting of a single bosonic chiral mode. The counting for the spin mode is more involved. We have
@@ Appendix F Different cylinder geometries
The transition to a semion crystal can only be observed on cylinder geometries, which are commensurate with the expected charge order of the crystal. We consider YCn-n and XCn-n cylinders, which have circumferences $L_{y}=n$. For YCn-n cylinders, we require $n=3\ell$ to be an integer multiple of $3$ and for XCn-n cylinders $n=2\ell$ has to be an even integer, to accommodate the CDW with every third site occupied. The CDW on YCn-n cylinders forms an effective XC(n1)-(n-1) geometry, while XCn-n reduces to YC(n2)-(n-2); see Fig. 8. For small cylinders, the details of the observed transition depend on the precise geometry, since we are effectively capturing a 1D transition. In the main text, we focused on the XC4 cylinder in Fig. 7. The equivalent data for the YC3 geometry is shown in Fig. 9.
Appendix G Computation of correlation lengths
In our tensor network simulations, Sec. IV, states are represented as infinite matrix product states (iMPS). In that case, the correlation function $C_{A}$ of any operator $A_{c}$ with charges $c$, can be expressed as [24]
$C_{A_{c}}(x)=\displaystyle\sum_{j}a_{j}\lambda_{c,j}^{x}.$ (84)
where $a_{j}$ are some constant factors and $\lambda_{c,j}$ is the $j$-th eigenvalue of the transfer matrix in charge sector $c$. For large $x\rightarrow\infty$, one finds
$C_{A_{c}}(x)\propto e^{-x/\xi_{c}}\quad\text{with}\quad\xi_{c}=-\frac{1}{\log\lambda_{c,1}}.$ (85)
Consequently, the correlations lengths $\xi_{c}$ are defined through the dominant eigenvalue of the transfer matrix in a given charge sector. In our iDMRG simulations, we have total charge $Q$, total spin $S_{z}$, and the momentum $k_{y}$ in $y$-direction, as conserved charges. The correlations length $\xi_{Q,S}=\max_{k_{y}}\xi_{Q,S,k_{y}}$ corresponds to the maximum (with respect to momentum $k_{y}$) correlation length for fixed charge $Q$ and spin $S_{z}$.
Appendix H Quantum critical point in terms of Cooper pairs
@@
where $\alpha$ and $\tilde{\alpha}$ are emergent gauge fields describing the bosonic hierarchy state, and $\varphi$ is a scalar that drives the crystallization transition. The plateau transition between the $2/3$ hierarchy state of Cooper pairs (corresponding to the fermionic FCI) and the $1/2$ parent state (corresponding to the semion crystal) is now described by a condensation transition of the $\varphi$ boson. If $\varphi$ is condensed (Mott) we obtain the $\nu^{\text{Cooper}}=2/3$ ($\nu^{\text{Cooper}}=1/2$) state of Cooper pairs. From the equations of motion we notice that the total electron density is given by $2\mathrm{d}\alpha/2\pi$, the density of $\varphi$ bosons is $\mathrm{d}\tilde{\alpha}/2\pi$ and the flux seen by the bosons $\mathrm{d}\beta=-3\mathrm{d}\alpha=-3\mathrm{d}\tilde{\alpha}/2$. The superconductor is then obtained when the bosons form a FQH state corresponding to their filling of 2/3-2/3.
Appendix I Generalization to doping the Jain sequence
While we focused on the spin-singlet Halperin sequence at filling $\nu=2/(2n\pm 1),\;n$ an even integer, discussed in the main text, our theoretical analysis can also be generalized to Jain sates of spinless fermions.
Transitions of Jain states.— Here we focus on the Jain sequence $\nu_{\text{Jain}}=\frac{p}{2p+1},p\in\mathbb{Z}$, which has a simple parton description in terms of
$e(x)=f_{1}(x)f_{2}(x)f_{3}(x),$ (87)
where the partons are all at integer fillings $\nu_{i}=\{p,1,1\}$ [60]. In contrast to the Halperin states, three fermionic partons suffice to describe the entire sequence. The Lagrangian of the Jain state after integrating out the constraints imposed by the parton construction can be expressed as
$\mathcal{L}_{\text{Jain}}=$- $\frac{2+\mathrm{sgn}(p)}{4\pi}\sum_{i=1}^{|p|}\alpha_{i}\wedge\mathrm{d}\alpha_{i}-\frac{2}{4\pi}\sum_{i\neq j}^{|p|}\alpha_{i}\wedge\mathrm{d}\alpha_{j}$ (88)
$+\frac{1}{2\pi}A\wedge\mathrm{d}\sum_{i=1}^{|p|}\alpha_{i}.$
To favor superconducting doped states, we again construct possible continuous transitions of these states to CDW’s by employing an extended parton construction
$e(x)=f_{1}f_{2}f_{3}b(x),$ (89)
where $b$ is a hard-core boson as before. When the boson is condensed we recover $\mathcal{L}_{\text{Jain}}$, while a crystallized boson leads to an insulating CDW with topological order phase, which can be thought of as an anyon crystal (AX) with action
$\mathcal{L}_{\text{AX}}=-\frac{2\mathrm{sgn}(p)}{4\pi}\sum_{i=1}^{|p|-1}\alpha_{i}\wedge\mathrm{d}\alpha_{i}-\frac{\mathrm{sgn}(p)}{4\pi}\sum_{i\neq j}^{|p|-1}\alpha_{i}\wedge\mathrm{d}\alpha_{j},$ (90)
with $|\mathrm{det}K_{\text{AX}}|=|p|$, leading to a residual topological order of the crystal. Note that the vortex which becomes soft has charge $Q_{\Phi}=e\nu=e\frac{p}{2p+1}$, which suggests a path to anyon-SC when $p\in 2\mathbb{Z}$. Perhaps even more exotic, when $p\in 2\mathbb{Z}+1$ the system could naturally form a charge-$p$ Fermi liquid.
Relation to plateau transition of composite objects.— As for the singlet states, stacking with integer phases highlights a dual description in terms of pleateau transitions of composite objects. Imagine stacking the critical theory with a $\nu=-p$ Chern insulator. The resulting insulator insulator transition involves a change of Hall conductivity
$\sigma_{xy}$ $:\quad\frac{p}{2p+1}\longleftrightarrow 0$ (91)
$\sigma^{\text{stacked}}_{xy}$ $:\quad\frac{-2p^{2}}{2p+1}\longleftrightarrow-p.$ (92)
Such a transition can be interpreted as a natural transition between a $\nu_{\text{composite}}=-1/p$ Laughlin state of charge-$p\cdot e$ particles and its hierarchy state at filling $\nu_{\text{composite}}=-2/(2p+1)$ describing the AX and the Jain state (the charge vector is $p$). As such the QCP is equivalent to the plateau transition between
$\nu_{\text{composite}}:\quad\frac{-2}{2p+1}\longleftrightarrow-\frac{1}{p}$ (93)
[…]
Based on the arguments presented in the previous sections we expect transitions of Jain states involving even $p$ to yield a charge $|p|$ superconductor.
Doping the Jain state.— For $p=\pm 2$ the physics is identical to the one discussed for the Halperin singlets, albeit in the absence of $\mathrm{SU(2)}$ symmetry. Nevertheless the resulting superconductor is found to have chiral central charge $c_{-}=\pm 2$.
The next interesting superconductor appears at $\nu=4/7$ (p=4p=-4). Repeating a derivation similar to the one in the main text, we find that the vortices are at filling $\nu_{\Phi}=7/4$ and have $7$ flavors. When translational invariance is spontaneously broken upon condensation, one of the vortices can naturally form a hierarchy state at $\nu=2-\frac{1}{4}$ around the bosonic integer quantum Hall state with $K$-matrix
$K=\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&-4\end{pmatrix},\qquad\mathbf{q}=\begin{pmatrix}1\\ 1\\ 1\end{pmatrix}$. (94)
Upon combining this with the fermionic sector, we indeed find that the resulting topological superconductor is charge-$4e$ with chiral central charge $c_{-}=-4$ and no residual topological order.
As in the previous cases, any other gapped vortex state at this filling would also yield a superconductor, although it could generically have topological order and in that case would have to be classified as an SC phase.